OpenStudy (anonymous):

cos(-7pi/12) Using half-angle I get: Sqrt((2-sqrt3)/4) Is this correct? What is the sign?

OpenStudy (psymon):

So you would be using -7pi/6, which is the same as 5pi/6 in the positive direction. Either way, I just like positive angles: \[\sqrt{\frac{ 1+\cos \frac{ 5 \pi }{ 6 } }{ 2 }}= \sqrt{\frac{ 1-\frac{ \sqrt{3} }{ 2 } }{ 2 }}\] \[\sqrt{\frac{ \frac{ 2-\sqrt{3} }{ 2 } }{ 2 }}= \sqrt{\frac{ 2-\sqrt{3} }{ 4 }}= \frac{ 1 }{ 2 }*\sqrt{2-\sqrt{3}}\] Look slike you did fine, you can just factor the 4 denominator out of the root.

OpenStudy (anonymous):

why isn't it negative?

OpenStudy (psymon):

The negative is inside of there. Theres nowhere else a negative could be really. No matter value of cosine any trig function you pick, youre going to have any negative sign inside of the root.

OpenStudy (anonymous):

i mean why isn't it -(1/2)sqrt(2-sqrt3) since cosine of -7pi/12 is negative

OpenStudy (psymon):

Youre right, I just left out the +/- in front of the formula as I was typing it.

OpenStudy (psymon):

I never put it in and only did the square root portion.

OpenStudy (anonymous):

hey thanks for the help! so its both +/- or only -?

OpenStudy (psymon):

Well, the formula should just naturally have a +/- in front of it. Then which one you select is knowledge of the quadrant youre in.

OpenStudy (anonymous):

k thanks